# An exact dimension-reduced dynamic theory for developable surfaces and curve-fold origami

### v1

@@ 4.1 Dynamics of a curve-folded plate
Figure 5: (a) FEM result of the dynamics of a curved plate. (b) Comparison between the theoretical results and FEM results.
A curve-folded plate is a basic cell of curve-fold origami. In this example, we analyse a freely developing curved-folded plate where the fold starts at the middle point along the length. According to (Fuchs and Tabachnikov, 1999), the generators of a freely curve-folded plate remain parallel, leading to the result that the panels can be simplified as a non-uniform Euler-Bernoulli beam. Specifically, the geometric variants are
$\eta_{\pm}=\pm\cot((S-\frac{S_{0}}{2})\kappa_{g}),$ (49)
$v_{0\pm}=\frac{L}{2}+\frac{1}{\kappa_{g}}(\cos(\frac{2S-S_{0}}{2\kappa_{g}^{2}})-\cos(\frac{S_{0}}{2\kappa_{g}^{2}})),$
$\gamma_{\pm}=\eta_{\pm}^{\prime}\mp(1+\eta_{\pm}^{2})\kappa_{g},$
and we assume the torsion remains zero. Equation (46) is thus simplified into
@@
$\mathbf{m}=-2(1+\eta^{4})Dv_{0}\eta\kappa_{n}\mathbf{e}_{1}-Dv_{0}\kappa_{n}(1-\eta^{2})\mathbf{e}_{2}+m_{3}\mathbf{e}_{3},$
@@
where $\eta_{\pm}$ and $v_{0\pm}$ are given in Eq. (49). Solving Eq. (50) yields the dynamic process of a developing curve-folded plate. The mechanical parameters are selected as $D\kappa_{g}/k_{c}=1$, and the geometric parameters are $L/W=1/2$. The process is also simulated in finite element analysis, showing good agreement with the theoretical results as demonstrated in Fig. 5.
@@ 4.3 Dynamics of a cut-then-fold ring origami
The equilibrium configurations of ring origami, in both closed and open forms, have been analyzed in Dias et al. (2012); Dias and Audoly (2014). These studies show that a folded closed ring buckles out of plane and acquires nonzero torsion along the crease, whereas a folded open ring remains planar with zero torsion. In our dynamic example, we examine the evolution of a cut folded closed ring. The resulting motion gradually drives the structure from an initially out-of-plane configuration toward a planar configuration, as illustrated in Fig. 7.
@@
Numerically solving Eq. (56) yields the curvature and torsion distributions at any time $t$. The theory indicates that the torsion of the cut ring gradually decreases while the curvature increases. This behavior is reproduced in finite-element simulations, as shown in Fig. 7, where the mechanical parameters satisfy $D\kappa_{g}/k_{c}=1$. A direct comparison between the theoretical and simulated curvature and torsion is also presented in Fig. 7, which demonstrates good quantitative agreement.
@@ Appendix B Component form of the dynamical equations in the fixed frame
$\quad\quad\quad\quad\quad\quad+\bm{\Omega}_{\pm}\times\left(\bm{\Omega}_{\pm}\times\left(\pm\eta_{\pm}\mathbf{t}\pm\cos\frac{\phi}{2}\mathbf{n}-\sin\frac{\phi}{2}\mathbf{b}\right)\right)\bigg)-\frac{1}{3}v_{0\pm}^{3}\dot{\eta}_{\pm}\bm{\Omega}_{\pm}^{\prime}\times\left(\pm\eta_{\pm}\mathbf{t}\pm\cos\frac{\phi}{2}\mathbf{n}-\sin\frac{\phi}{2}\mathbf{b}\right)\bigg],$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\frac{1}{4}v_{0\pm}^{4}\dot{\eta}_{\pm}\left(\pm\eta_{\pm}\mathbf{t}\pm\cos\frac{\phi}{2}\mathbf{n}-\sin\frac{\phi}{2}\mathbf{b}\right)\times\left(\bm{\Omega}_{\pm}^{\prime}\times\left(\pm\eta_{\pm}\mathbf{t}\pm\cos\frac{\phi}{2}\mathbf{n}-\sin\frac{\phi}{2}\mathbf{b}\right)\right)\bigg],$
$\mathbf{m}_{\mp}=\pm\frac{1}{\kappa_{n}}\left(-\frac{\partial}{\partial S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}+\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}}\right)\mathbf{t}+\left(\pm m_{3}\sin\frac{\phi}{2}\mp\left(\frac{\partial\varepsilon_{\pm}}{\partial\kappa_{n}}+\frac{\eta_{\pm}}{\kappa_{n\pm}}(\frac{\partial}{\partial S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}-\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}})\right)\cos\frac{\phi}{2}\right)\mathbf{n}$
$\quad\quad+\left(m_{3}\cos\frac{\phi}{2}+\left(\frac{\partial\varepsilon_{\pm}}{\partial\kappa_{n}}+\frac{\eta_{\pm}}{\kappa_{n\pm}}(\frac{\partial}{\partial S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}-\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}})\right)\sin\frac{\phi}{2}\right)\mathbf{b},$
@@
According to Eq. (44), the angular velocity of the Darboux frame $\mathbf{\Omega}_{\pm}$ relates to $\mathbf{\Omega}_{f}$, following
$\mathbf{\Omega}_{\pm}=\mathbf{\Omega}_{f}\pm\frac{\dot{\phi}}{2}\mathbf{t}.$ (B.11)
@@
Substituting Eqs. (44), (B.8)-(B.11) into (B.7) yields 7 equations related to $x,y,z,\psi,\theta,\varphi,F_{x\Sigma},F_{y\Sigma},F_{z\Sigma},m_{3\pm}$. Along with Eq. (B.12) and $\psi^{\prime}\sin\theta\cos\varphi-\theta^{\prime}\sin\varphi=0$, we derive 11 equations with 11 variables, which can be solved numerically in a closed form.

### v2

@@ 4.1 Dynamics of a curve-folded plate
Figure 5: (a) FEM result of the dynamics of a curve-folded plate. (b) Comparison between the theoretical results and FEM results.
A curve-folded plate is a basic cell of curve-fold origami. In this example, we analyse a freely developing curved-folded plate where the fold starts at the middle point along the length. The specific loading process is to press the two long edges to fold the plate statically, and then release the compression to unfold it dynamically. Previous studies indicate that the generators remain parallel in the static deformation (Lee et al., 2018). Here, we first assume that this conclusion remains applicable in dynamics and will verify it subsequently. Under such an assumption, the geometric variants are
$\eta_{\pm}=\tan(\frac{2S-S_{0}}{2}\kappa_{g}),$ (50)
$v_{0\pm}=\cos(\frac{2S-S_{0}}{2}\kappa_{g})\frac{L}{2}\mp\frac{\cos(\frac{2S-S_{0}}{2}\kappa_{g})}{\kappa_{g}}(\cos(\frac{2S-S_{0}}{2}\kappa_{g})-\cos(\frac{S_{0}}{2}\kappa_{g})),$
$\gamma_{\pm}=\pm\eta_{\pm}^{\prime}\mp(1+\eta_{\pm}^{2})\kappa_{g}=0.$
Eq. (47) is thus simplified into
@@
$\mathbf{M}_{\pm}=Dv_{0}(\mp 2(1+\eta_{\pm}^{4})\eta_{\pm}\kappa_{n\pm}+\frac{1}{2}v_{0}((1+\eta_{\pm}^{2})\kappa_{n\pm})^{\prime})\mathbf{e}_{1}-Dv_{0}(\pm\kappa_{n\pm}(1-\eta_{\pm}^{2})+\frac{\eta_{\pm}}{2}v_{0}((1+\eta_{\pm}^{2})\kappa_{n\pm})^{\prime})\mathbf{e}_{2}+m_{3}\mathbf{e}_{3},$
@@
where $\eta_{\pm}$ and $v_{0\pm}$ are given in Eq. (50). Solving Eq. (51) yields $\mathbf{\Omega}_{-}\times\mathbf{l-}=0$, indicating that the solution does not violate the assumption that the generators remain parallel and the panels deform cylindrically. Since the dynamic process is uniquely determined by Eq. (47), the solution of Eq. (51) is the unique solution that satisfies Eq. (32) in this example. The process is also simulated in finite element analysis, where the mechanical parameters are selected as $D\kappa_{g}/k_{c}=1$, and the geometric parameters are $L/W=1/2$. The good agreement between the theoretical and numerical results are demonstrated in Fig. 5.
@@ 4.3 Dynamics of a cut-then-fold ring origami
The equilibrium configurations of ring origami, in both closed and open forms, have been analyzed in Dias et al. (2012); Dias and Audoly (2014). These studies show that a folded closed ring buckles out of plane and acquires nonzero torsion along the crease, whereas a folded open ring remains planar with zero torsion. In our dynamic example, we examine the evolution of a cut folded closed ring. The resulting motion gradually drives the structure from an initial out-of-plane configuration to a planar configuration, as illustrated in Fig. 7.
@@
Numerically solving Eq. (57) yields the curvature and torsion distributions at any time $t$. The theory indicates that the torsion of the cut ring gradually decreases while the curvature increases. This behavior is reproduced in finite-element simulations, as shown in Fig. 7, where the mechanical parameters satisfy $D\kappa_{g}/k_{c}=1$ and $\phi_{0}=\pi$. A direct comparison between the theoretical and simulated curvature and torsion is also presented in Fig. 7, which demonstrates good quantitative agreement.
@@ Appendix B Component form of the dynamical equations in the fixed frame
$\quad\quad\quad\quad\quad\quad+\bm{\Omega}_{\pm}\times\left(\bm{\Omega}_{\pm}\times\left(\pm\eta_{\pm}\mathbf{t}\pm\cos\frac{\phi}{2}\mathbf{n}-\sin\frac{\phi}{2}\mathbf{b}\right)\right)\bigg)\mp\frac{1}{3}v_{0\pm}^{3}\dot{\eta}_{\pm}\bm{\Omega}_{\pm}^{\prime}\times\left(\pm\eta_{\pm}\mathbf{t}\pm\cos\frac{\phi}{2}\mathbf{n}-\sin\frac{\phi}{2}\mathbf{b}\right)\bigg],$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\mp\frac{1}{4}v_{0\pm}^{4}\dot{\eta}_{\pm}\left(\pm\eta_{\pm}\mathbf{t}\pm\cos\frac{\phi}{2}\mathbf{n}-\sin\frac{\phi}{2}\mathbf{b}\right)\times\left(\bm{\Omega}_{\pm}^{\prime}\times\left(\pm\eta_{\pm}\mathbf{t}\pm\cos\frac{\phi}{2}\mathbf{n}-\sin\frac{\phi}{2}\mathbf{b}\right)\right)\bigg],$
$\mathbf{M}_{\pm}=\frac{1}{\kappa_{n\pm}}\left(-\frac{\partial}{\partial S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}+\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}}\right)\mathbf{t}+\left(m_{3}\sin\frac{\phi}{2}-\left(\frac{\partial\varepsilon_{\pm}}{\partial\kappa_{n\pm}}+\frac{\eta_{\pm}}{\kappa_{n\pm}}(\frac{\partial}{\partial S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}-\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}})\right)\cos\frac{\phi}{2}\right)\mathbf{n}$
$\quad\quad\pm\left(m_{3}\cos\frac{\phi}{2}+\left(\frac{\partial\varepsilon_{\pm}}{\partial\kappa_{n\pm}}+\frac{\eta_{\pm}}{\kappa_{n\pm}}(\frac{\partial}{\partial S}\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}^{\prime}}-\frac{\partial\varepsilon_{\pm}}{\partial\eta_{\pm}})\right)\sin\frac{\phi}{2}\right)\mathbf{b},$
@@
Substituting Eqs. (44),(45) (B.8)-(B.10) into (B.7) yields 7 equations related to $x,y,z,\psi,\theta,\varphi,F_{x\Sigma},F_{y\Sigma},F_{z\Sigma},m_{3\pm}$. Along with Eq. (B.11) and $\psi^{\prime}\sin\theta\cos\varphi-\theta^{\prime}\sin\varphi=0$, we derive 11 equations with 11 variables, which can be solved numerically in a closed form.